OpenStudy (anonymous):

You spin the spinner below once. What is the probability that the arrow will land on an even number? A. 1/8 B. 3/8 C. 1/4 D. 1/2

OpenStudy (anonymous):

@iGreen would it be 1/4?

OpenStudy (igreen):

No..

OpenStudy (igreen):

Count the even numbers on there.

OpenStudy (igreen):

Then count the total amount of spaces.

OpenStudy (anonymous):

so 4 and 8, then what?

OpenStudy (igreen):

That's our fraction. 4/8 Now simplify it

OpenStudy (anonymous):

4/8

OpenStudy (anonymous):

2/4

OpenStudy (anonymous):

2/4 and then simplify it again?

OpenStudy (solomonzelman):

A probability that it will lend on a specific number (lets say, only on number "3") is 1/8. Because it can lend on 1 fielf out of 8. Now, for it to lend on any number 1-7 (including 1 and 7) is 7/8 because it can lend on 7 possible fields out of all 8. When you want it to lend on an even number: 1) how many fields with even numbers there are? 2) How many possibilities is it out of 8? 3) What is (therefore) your answer ?

OpenStudy (anonymous):

1/2

OpenStudy (igreen):

Yes @vicyoria15

OpenStudy (anonymous):

So D.

OpenStudy (igreen):

@xBlackDimondx Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. http://openstudy.com/code-of-conduct

OpenStudy (solomonzelman):

yes, 1.2 is the correct answer.

OpenStudy (solomonzelman):

I meant 1/2

OpenStudy (anonymous):

Thanks to all of you!!

OpenStudy (anonymous):

Because there are 4 even numbers in the number 8

OpenStudy (anonymous):

So its 4/8 = 2/4=1/2

OpenStudy (anonymous):

Your Welcome!

OpenStudy (anonymous):

A bag contains 5 black marbles, 9 white marbles, and 2 red marbles. One marble is to be chosen at random. What is the probability that the chosen marble will be red? A. 1/16 B. 1/8 C. 1/3 D. 1/2